# Modeling spectral filtering effects on color-matching functions: Implications for observer variability

### v1

@@
This study investigates the impact of spectral filtering on color-matching functions (CMFs) and its implications for observer variability modeling. We conducted color matching experiments with a single observer, both with and without a spectral filter in front of a bipartite field. Using a novel computational approach, we estimated the filter transmittance and transformation matrix necessary to convert unfiltered CMFs to filtered CMFs. Statistical analysis revealed good agreement between estimated and measured filter characteristics, particularly in central wavelength regions. Applying this methodology to compare between Stiles and Burch 1955 (SB1955) mean observer CMFs and our previously published "ICVIO" mean observer CMFs, we identified a "yellow" (short-wavelength suppressing) filter that effectively transforms between these datasets. This finding aligns with our hypothesis that observed differences between the CMF sets are attributable to age-related lens yellowing (average observer age: 49 years in ICVIO versus 30 years in SB1955). Our approach enables efficient representation of observer variability through a single filter rather than three separate functions, offering potentially reduced experimental overhead while maintaining accuracy in characterizing individual color vision differences.
@@ 1 Introduction
Accurate color quantification is a cornerstone of color science, underpinning diverse applications such as colorimetry, color reproduction, material appearance perception, and color management. Central to this quantification are color-matching functions (CMFs), which describe the amounts of three reference stimuli required to match monochromatic test stimuli across the visible spectrum [1].
@@
The Commission Internationale de l’Éclairage (CIE) has standardized CMFs by aggregating experimental data to represent the average color-matching responses of individuals with normal color vision [2]. However, discrepancies between the predictions of standard CMFs and actual experimental matches made by color-normal observers have been widely documented [3, 4, 5, 6]. These deviations highlight an essential feature of human color vision: individual variability.
@@
Such variability arises from physiological differences in the human visual system, including variations in lens optical density [7], macular pigment density [8], genotypic differences in cone photopigments [9, 10], and age-related changes in ocular media [11]. Understanding these individual differences has significant implications for industries reliant on precise color reproduction.
@@
Efforts to characterize inter-observer variability in color-matching functions (CMFs) have taken different approaches. Asano et al. [5, 6] developed categorical models to classify observers based on their color-matching behavior, providing a framework for quantifying variability across individuals. Other studies, such as those by Shi et al. [12], have aimed to estimate physiological parameters—such as cone spectral sensitivities—that best fit an observer’s color matches, revealing deviations from standardized models like the CIE 2006 observer [13]. These approaches have advanced our understanding of inter-observer differences but primarily focus on modeling variability rather than exploring its underlying causes.
@@
Andersen et al. [14] tackled a different problem: estimating individual cone fundamentals from CMFs by combining lens and macular pigment effects into a single "pre-filter". This approach provided a way to relate individual CMFs to cone fundamentals through a linear optimization framework, offering insights into how optical filtering contributes to inter-observer variability. Inspired by this concept, our study aims to emulate the effect of Andersen’s "pre-filter" experimentally by introducing a physical spectral filter into a color-matching setup.
@@
Recent advances in computational filter design for imaging systems provide a framework for modeling transformations induced by spectral filters. Finlayson and Zhu [15, 16, 17] developed methods to optimize spectral filters that transform camera spectral sensitivities to approximate the CIE XYZ color-matching functions. These methods employ alternating least-squares (ALS) optimization to minimize colorimetric errors while meeting physical constraints such as non-negative transmittance and smoothness. Integrating these filters with linear transformations has demonstrated significant improvements in imaging system accuracy [18].
@@
A key concept from these works is using pre-filters to modify a device’s spectral characteristics, aligning them with standardized color representations. For instance, Finlayson and Zhu [15] showed that an optimized filter placed in front of a camera can achieve near-perfect compliance with the Luther condition—a criterion ensuring that camera sensitivities are linear combinations of human CMFs. This principle illustrates how spectral filtering bridges device-specific responses and standardized models.
@@ 2 Current Study
This study investigates the impact of an spectral filter on color matching behavior for a single observer. By maintaining consistent experimental conditions and varying only the filter’s presence in front of the bipartite field, we aim to isolate its effect on color matching. This approach enables direct comparison of color-matching functions (CMFs) obtained with and without the filter, providing insights into the filter’s influence on the observer’s color matching process and potential reflections of inherent perceptual variations.
@@ 2.1 Apparatus and Setup
The experimental setup and method employed in this study are based on the apparatus and procedures detailed in our previous work [19], with a slight modification to accommodate the current experiment. The core component of the experimental apparatus is a custom-built colorimeter designed to measure individual color-matching functions (CMFs) through psychophysical experiments. This apparatus, constructed using 3D-printed integrating chambers, optical elements, and RGB LED light engines, creates a bipartite field allowing the observer to view two juxtaposed fields of light.
@@
The "purple" filter was specifically chosen because it attenuates the edges of the visible spectrum minimally, which is particularly advantageous for this experiment. At these edges, we already observe low luminance levels, as established in our previous work [19]. By minimizing attenuation in these regions, the filter ensures that the inherent low luminance at spectral boundaries does not interfere with the color-matching process. This makes it highly compatible with the existing experimental setup, allowing for precise and reliable comparisons between matches made with and without the filter in place.
@@ 2.2 Experimental Method
The experiment was conducted in a dark room, with the bipartite field serving as the only light source. The observer’s eyes were positioned approximately 80 cm from the bipartite field, corresponding to a visual angle of about 1.4°. This ensured that small head movements did not significantly impact color matching. Therefore, no head restraints were used, but the observer was instructed to loosely maintain the same viewing position. The experiment began with a near-spectral test light of peak wavelength $\lambda_{i}$ set on the test side of the bipartite field, while the RGB LED lights, which in this experiment serve as reference lights, were initially off. The observer was instructed to freely scan the bipartite field with both eyes and adjust the intensities of the RGB reference lights to match the color of the reference field to that of the test field. In so doing, they were instructed to minimize differences in hue, brightness, and saturation between the two fields. Once a match was achieved, the SPDs of the test field, $\mathbf{T}_{\lambda_{i}}$ and the reference field, $\mathbf{M}_{\lambda_{i}}$ were measured using a spectroradiometer. Then, the reference lights were reset to zero, and a third measurement of the test light was taken by itself, $\mathbf{S}_{\lambda_{i}}$. This procedure was repeated for test lights with peak wavelengths ranging from 400 nm to 720 nm at every 10 nm interval. Several measurements (at least three, c.f. [19] for details) were made for each test light, to account for the observer’s matching uncertainty. This experiment was repeated with the filter placed between the observer’s eyes and the bipartite field, effectively integrating it as a transient component of the visual system during color matching. However, it is critical to note that only the perceptual matching process utilized the filter. Upon achieving a match, the filter was removed prior to any SPD measurements to ensure all radiometric data reflected the physical stimuli without spectral filtering. This approach isolates the filter’s perceptual influence while preserving consistency in radiometric measurements for direct comparison between conditions.
@@ 2.3.1 Computation of color-matching coefficients
Two sets of raw color-matching data were obtained from a single observer. One set corresponds to the experiment conducted without any filter, while the other set represents the case where a filter was placed in front of the bipartite field. The computation of tristimulus values followed the methodology outlined in our previous work [19]. The process involved:
@@
The resulting tristimulus values, expressed as functions of wavelength, are presented in Fig. 2 for both experimental conditions (with and without filter). This figure illustrates the tristimulus values relative to unit-energy normalized reference color stimuli R, G, and B, with peak wavelengths of 634 nm, 534 nm, and 452 nm, respectively. At each test wavelength, multiple data points are plotted, corresponding to the number of color matches made at that wavelength. This representation allows for a direct comparison of the filter’s effect on the observer’s color-matching behavior across the visible spectrum.
@@
Figure 2: Color-matching coefficients from color-matching experiment done : (a) without filter and (b) with filter
@@ 2.3.2 Filter Estimation Algorithm
This optimisation problem can be solved using quadratic programming or Lagrangian multipliers [18]. The two terms in the optimization are both important. The x minimizing $|\textbf{A}\textbf{x}|^{2}$ alone would be in the null-space of $A$, the dimension of which is 3N9N=2N93N-9-N=2N-9 (there would be many—in fact a whole space of—solutions). A particular solution can be found thanks to the smoothness constraint $\alpha\left|\tilde{\textbf{T}}\textbf{x}\right|^{2}$.
@@ 3.1 Mean color-matching functions (CMFs)
The mean CMFs for the observer under both experimental conditions, with and without the "purple" filter, are illustrated in Fig. 3. The CMFs are plotted as functions of wavelength, with error bars representing the 95 % bootstrapped confidence intervals (CI) of the mean. This figure provides a visual comparison of how the presence of the filter affects the observer’s color-matching behavior across the visible spectrum.
@@
Figure 3: Mean CMFs with error bars representing 95% CI of the mean.
@@ 3.2 Filter estimation and uncertainty analysis
Fig. 4 displays the first 20 of N (where N = 1000) filter estimates derived from the bootstrapped filter estimation process. This visualization highlights the variability in the estimated filter characteristics, showcasing the range of possible filter transmittance vectors, $\hat{\textbf{f}}^{(i)}$, that could transform the CMFs without the filter to those with the filter, within the bounds of experimental noise.
Figure 4: The first 20 of N filter estimates
The mean estimated filter, along with error bars representing two times the standard deviation ($2\sigma_{\mathbf{f}}$) of the filter estimate at each wavelength from 400 nm to 720 nm, at 10 nm intervals, is presented in Figure 5. This figure allows for a direct comparison with the ground truth filter SPD measurement, providing insight into the accuracy and reliability of the filter estimation algorithm. It is worth noting that the estimated filter transmittance is normalized to have a sum of 1, as only the shape of the filter’s transmittance curve can be estimated, not the absolute transmittance. Similarly, the resampled ground truth of the filter transmittance is also normalized in this way to facilitate comparison.
Figure 5: Mean estimated filter and the resampled ground truth of the filter, error bars representing 2 $\sigma_{\mathbf{f}}(\lambda)$
@@ 4.1 Performance of Filter estimation algorithm
The experimental results validate the hypothesis that a spectral filter (and a linear transformation matrix) can be computed to convert unfiltered CMFs to those obtained with a filter. The estimated filter’s spectral transmittance closely aligns, at least in the central region, with the measured transmission properties of the physical filter used in the experiment. This alignment underscores the robustness of the filter estimation algorithm in capturing the spectral characteristics of physical filters used in color-matching experiments. It can be observed that under 440 nm and above 690 nm, the estimated filter is no longer in agreement with the ground truth. This is potentially due to reduced illumination induced by filtering during the color-matching experiments. The observer did report that matching is harder at the edges of the visible spectrum under the filtered experimental condition, which is also evident given the larger error bars of the estimated filter in those regions.
@@ 4.2 Transformation between two different CMFs sets
Given the promising results of the filter estimation algorithm, we decided to compute the spectral filter and $3\times 3$ matrix required to transform between the mean CMFs from two different datasets. The CMFs chosen for this exercise are the Stiles and Burch 1955 2 mean observer CMFs, and the 2 mean observer CMFs from our previously published work[19], which we are calling the ICVIO mean observer CMFs here. The two CMFs sets are plotted in Fig. 6.
@@
The filter that would convert from the Stiles and Burch 1955 2 mean observer CMFs to that of the ICVIO mean observer CMFs is shown in Fig. 7. The computed filter exhibits a distinctive spectral profile, with progressively decreasing transmittance toward shorter wavelengths and relatively higher transmittance in the middle and long wavelength regions. This spectral shape closely resembles that of a "yellow" filter, with significant attenuation in the blue region of the spectrum. This finding is particularly noteworthy as it aligns with well-documented physiological changes that occur in the human visual system with age—specifically, the yellowing of the crystalline lens[7]. The lens naturally accumulates yellow pigmentation over time, which selectively absorbs short-wavelength light and alters color perception, particularly in the blue region of the spectrum. This result was expected, given that the mean age of the observers comprising the ICVIO dataset is 49 years old, compared to mean age of 30 years old for observers in the Stiles and Burch 1955 dataset. This adds weight to our previous hypothesis [19] that the discrepancy observed between the Stiles and Burch 1955 2 mean observer CMFs and our own ICVIO mean observer CMFs, can be attributed, at least partially, to physiological factors such as lens aging.

### v2

@@
This study investigates the impact of spectral filtering on color-matching functions (CMFs) and its implications for observer variability modeling. We conducted color matching experiments with two observers, both with and without a spectral filter in front of a bipartite field. Using a novel computational approach, we estimated the filter transmittance and transformation matrix necessary to convert unfiltered CMFs to filtered CMFs. Statistical analysis revealed good agreement between estimated and measured filter characteristics, particularly in central wavelength regions. Applying this methodology to compare between Stiles and Burch 1955 (SB1955) mean observer CMFs and our previously published "ICVIO" mean observer CMFs, we identified a "yellow" (short-wavelength suppressing) filter that effectively transforms between these datasets. This finding aligns with our hypothesis that observed differences between the CMF sets are attributable to age-related lens yellowing (average observer age: 49 years in ICVIO versus 30 years in SB1955). Our approach enables efficient representation of observer variability through a single filter rather than three separate functions, offering potentially reduced experimental overhead while maintaining accuracy in characterizing individual color vision differences.
@@ 1 Introduction
Accurate color quantification is a cornerstone of color science, underpinning diverse applications such as colorimetry, color reproduction, material appearance perception, and color management. Central to this quantification are color-matching functions (CMFs), which describe the amounts of three reference stimuli required to match monochromatic test stimuli across the visible spectrum [Ohta_Robertson2005colorimetry].
@@
The Commission Internationale de l’Éclairage (CIE) has standardized CMFs by aggregating experimental data to represent the average color-matching responses of individuals with normal color vision [Schanda2007CIEColorimetry]. However, discrepancies between the predictions of standard CMFs and actual experimental matches made by color-normal observers have been widely documented [csuti2008colour, sarkar2011identification, asano2015individual, asano2016color]. These deviations highlight an essential feature of human color vision: individual variability.
@@
Such variability arises from physiological differences in the human visual system, including variations in lens optical density [pokorny1987aging], macular pigment density [smith1976variability], genotypic differences in cone photopigments [thomas2011effect, stockman200042], and age-related changes in ocular media [artigas2012spectral]. Understanding these individual differences has significant implications for industries reliant on precise color reproduction.
@@
Efforts to characterize inter-observer variability in color-matching functions (CMFs) have taken different approaches. Asano et al. [asano2015individual, asano2016color] developed categorical models to classify observers based on their color-matching behavior, providing a framework for quantifying variability across individuals. Other studies, such as those by Shi et al. [shi2024multi], have aimed to estimate physiological parameters—such as cone spectral sensitivities—that best fit an observer’s color matches, revealing deviations from standardized models like the CIE 2006 observer [cie2006fundamental]. These approaches have advanced our understanding of inter-observer differences but primarily focus on modeling variability rather than exploring its underlying causes.
@@
Andersen et al. [andersen2016estimating] tackled a different problem: estimating individual cone fundamentals from CMFs by combining lens and macular pigment effects into a single "pre-filter". This approach provided a way to relate individual CMFs to cone fundamentals through a linear optimization framework, offering insights into how optical filtering contributes to inter-observer variability. Inspired by this concept, our study aims to emulate the effect of Andersen’s "pre-filter" experimentally by introducing a physical spectral filter into a color-matching setup.
@@
Recent advances in computational filter design for imaging systems provide a framework for modeling transformations induced by spectral filters. Finlayson and Zhu [finlayson2019finding, FinlaysonZhuVora, FinlaysonZhu] developed methods to optimize spectral filters that transform camera spectral sensitivities to approximate the CIE XYZ color-matching functions. These methods employ alternating least-squares (ALS) optimization to minimize colorimetric errors while meeting physical constraints such as non-negative transmittance and smoothness. Integrating these filters with linear transformations has demonstrated significant improvements in imaging system accuracy [andersen2025].
@@
A key concept from these works is using pre-filters to modify a device’s spectral characteristics, aligning them with standardized color representations. For instance, Finlayson and Zhu [finlayson2019finding] showed that an optimized filter placed in front of a camera can achieve near-perfect compliance with the Luther condition—a criterion ensuring that camera sensitivities are linear combinations of human CMFs. This principle illustrates how spectral filtering bridges device-specific responses and standardized models.
@@ 2 Current Study
This study investigates the impact of an spectral filter on color matching behavior for two observers, Observer-I (male, age 52) and Observer-C (female, age 30). By maintaining consistent experimental conditions and varying only the filter’s presence in front of the bipartite field, we aim to isolate its effect on color matching. This approach enables direct comparison of color-matching functions (CMFs) obtained with and without the filter, providing insights into the filter’s influence on the observers’ color matching process and potential reflections of inherent perceptual variations.
@@ 2.1 Apparatus and Setup
The experimental setup and method employed in this study are based on the apparatus and procedures detailed in our previous work [ragoo2024apparatus], with a slight modification to accommodate the current experiment. The core component of the experimental apparatus is a custom-built colorimeter designed to measure individual color-matching functions (CMFs) through psychophysical experiments. This apparatus, constructed using 3D-printed integrating chambers, optical elements, and RGB LED light engines, creates a bipartite field allowing the observer to view two juxtaposed fields of light.
@@
The "purple" filter was specifically chosen because it attenuates the edges of the visible spectrum minimally, which is particularly advantageous for this experiment. At these edges, we already observe low luminance levels, as established in our previous work [ragoo2024apparatus]. By minimizing attenuation in these regions, the filter ensures that the inherent low luminance at spectral boundaries does not interfere with the color-matching process. This makes it highly compatible with the existing experimental setup, allowing for precise and reliable comparisons between matches made with and without the filter in place.
@@ 2.2 Experimental Method
The experiment was conducted in a dark room, with the bipartite field serving as the only light source. The observer’s eyes were positioned approximately 80 cm from the bipartite field, corresponding to a visual angle of about 1.4°. This ensured that small head movements did not significantly impact color matching. Therefore, no head restraints were used, but the observer was instructed to loosely maintain the same viewing position. The experiment began with a near-spectral test light of peak wavelength $\lambda_{i}$ set on the test side of the bipartite field, while the RGB LED lights, which in this experiment serve as reference lights, were initially off. The observer was instructed to freely scan the bipartite field with both eyes and adjust the intensities of the RGB reference lights to match the color of the reference field to that of the test field. In so doing, they were instructed to minimize differences in hue, brightness, and saturation between the two fields. Once a match was achieved, the SPDs of the test field, $\mathbf{T}_{\lambda_{i}}$ and the reference field, $\mathbf{M}_{\lambda_{i}}$ were measured using a spectroradiometer. Then, the reference lights were reset to zero, and a third measurement of the test light was taken by itself, $\mathbf{S}_{\lambda_{i}}$. This procedure was repeated for test lights with peak wavelengths ranging from 400 nm to 720 nm at every 10 nm interval. Several measurements (at least three, c.f. [ragoo2024apparatus] for details) were made for each test light, to account for the observer’s matching uncertainty. This experiment was repeated with the filter placed between the observer’s eyes and the bipartite field, effectively integrating it as a transient component of the visual system during color matching. However, it is critical to note that only the perceptual matching process utilized the filter. Upon achieving a match, the filter was removed prior to any SPD measurements to ensure all radiometric data reflected the physical stimuli without spectral filtering. This approach isolates the filter’s perceptual influence while preserving consistency in radiometric measurements for direct comparison between conditions.
@@ 2.3.1 Computation of color-matching coefficients
For each observer, two sets of raw color-matching data were obtained. One set corresponds to the experiment conducted without any filter, while the other set represents the case where a filter was placed in front of the bipartite field. The computation of tristimulus values followed the methodology outlined in our previous work [ragoo2024apparatus]. The process involved:
@@
The resulting tristimulus values (for Observer-I), expressed as functions of wavelength, are presented in Fig. 2 for both experimental conditions (with and without filter). This figure illustrates the tristimulus values relative to unit-energy normalized reference color stimuli R, G, and B, with peak wavelengths of 634 nm, 534 nm, and 452 nm, respectively. At each test wavelength, multiple data points are plotted, corresponding to the number of color matches made at that wavelength. This representation allows for a direct comparison of the filter’s effect on the observer’s color-matching behavior across the visible spectrum.
@@
Figure 2: Color-matching coefficients from color-matching experiment done (for Observer-I) : (a) without filter and (b) with filter
@@ 2.3.2 Filter Estimation Algorithm
This optimisation problem can be solved using quadratic programming or Lagrangian multipliers [andersen2025]. The two terms in the optimization are both important. The x minimizing $|\textbf{A}\textbf{x}|^{2}$ alone would be in the null-space of $A$, the dimension of which is 3N9N=2N93N-9-N=2N-9 (there would be many—in fact a whole space of—solutions). A particular solution can be found thanks to the smoothness constraint $\alpha\left|\tilde{\textbf{T}}\textbf{x}\right|^{2}$.
@@ 3.1 Mean color-matching functions (CMFs)
The mean CMFs for the two observers under both experimental conditions, with and without the "purple" filter, are illustrated in Fig. 3 and Fig. 4. The CMFs are plotted as functions of wavelength, with error bars representing the 95 % bootstrapped confidence intervals (CI) of the mean. This figure provides a visual comparison of how the presence of the filter affects each observer’s color-matching behavior across the visible spectrum.
@@
Figure 3: Observer-I : Mean CMFs with error bars representing 95% CI of the mean.
(a) Experimental condition : no filter
(b) Experimental condition : with filter
Figure 4: Observer-C : Mean CMFs with error bars representing 95% CI of the mean.
@@ 3.2 Filter estimation and uncertainty analysis
Fig. 5 displays the first 20 of N (where N = 1000) filter estimates (for Observer-I) derived from the bootstrapped filter estimation process. This visualization highlights the variability in the estimated filter characteristics, showcasing the range of possible filter transmittance vectors, $\hat{\textbf{f}}^{(i)}$, that could transform the CMFs without the filter to those with the filter, within the bounds of experimental noise.
Figure 5: The first 20 of N filter estimates (from Observer-I’s data)
The mean estimated filter, along with error bars representing two times the standard deviation ($2\sigma_{\mathbf{f}}$) of the filter estimate at each wavelength from 400 nm to 720 nm, at 10 nm intervals, is presented in Fig. 6 for Observer-I and in Fig. 7 for Observer-C. This figure allows for a direct comparison with the ground truth filter SPD measurement, providing insight into the accuracy and reliability of the filter estimation algorithm. It is worth noting that each estimated filter transmittance is normalized to have a sum of 1, as only the shape of the filter’s transmittance curve can be estimated, not the absolute transmittance. Similarly, the resampled ground truth of the "purple" filter transmittance is also normalized in this way to facilitate comparison.
Figure 6: Observer-I : Mean estimated filter and the resampled ground truth of the filter, error bars representing 2 $\sigma_{\mathbf{f}}(\lambda)$
Figure 7: Observer-C : Mean estimated filter and the resampled ground truth of the filter, error bars representing 2 $\sigma_{\mathbf{f}}(\lambda)$
@@ 4.1 Performance of Filter estimation algorithm
The experimental results validate the hypothesis that a spectral filter (and a linear transformation matrix) can be computed to convert unfiltered CMFs to those obtained with a filter. The estimated filter’s spectral transmittance closely aligns, at least in the central region, with the measured transmission properties of the physical filter used in the experiment, for each observer. This alignment underscores the robustness of the filter estimation algorithm in capturing the spectral characteristics of physical filters used in color-matching experiments. It can be observed that under 440 nm and above 690 nm, the estimated filters is no longer in agreement with the ground truth. This is potentially due to reduced illumination induced by filtering during the color-matching experiments. The observers did report that matching is harder at the edges of the visible spectrum under the filtered experimental condition, which is also evident given the larger error bars of the estimated filter in those regions.
@@ 4.2 Transformation between two different CMFs sets
Given the promising results of the filter estimation algorithm, we decided to compute the spectral filter and $3\times 3$ matrix required to transform between the mean CMFs from two different datasets. The CMFs chosen for this exercise are the Stiles and Burch 1955 2 mean observer CMFs, and the 2 mean observer CMFs from our previously published work[ragoo2024apparatus], which we are calling the ICVIO mean observer CMFs here. The two CMFs sets are plotted in Fig. 8.
@@
The filter that would convert from the Stiles and Burch 1955 2 mean observer CMFs to that of the ICVIO mean observer CMFs is shown in Fig. 9. The computed filter exhibits a distinctive spectral profile, with progressively decreasing transmittance toward shorter wavelengths and relatively higher transmittance in the middle and long wavelength regions. This spectral shape closely resembles that of a "yellow" filter, with significant attenuation in the blue region of the spectrum. This finding is particularly noteworthy as it aligns with well-documented physiological changes that occur in the human visual system with age—specifically, the yellowing of the crystalline lens[pokorny1987aging]. The lens naturally accumulates yellow pigmentation over time, which selectively absorbs short-wavelength light and alters color perception, particularly in the blue region of the spectrum. This result was expected, given that the mean age of the observers comprising the ICVIO dataset is 49 years old, compared to mean age of 30 years old for observers in the Stiles and Burch 1955 dataset. This adds weight to our previous hypothesis [ragoo2024apparatus] that the discrepancy observed between the Stiles and Burch 1955 2 mean observer CMFs and our own ICVIO mean observer CMFs, can be attributed, at least partially, to physiological factors such as lens aging.